Leisure Lodge Corporation is expected to pay the following dividends over the next four years: $22.00, $10.00, $8.20 and $2.80. Afterwards, the company pledges to maintain a constant 5 percent growth rate in dividends forever. If the required return on the stock is 16 percent, what is the current share price? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Share price
Answer #1

Value after year 4=(D4*Growth rate)/(Required return-Growth rate)

=(2.8*1.05)/(0.16-0.05)=$26.72727(APPROX)

Hence current share price=Future dividends*Present value of discounting factor(16%,time period)

=22/1.16+10/1.16^2+8.20/1.16^3+2.8/1.16^4+26.72727/1.16^4

which is equal to

=$47.96(Approx).